OpenStudy (zzelinski):

0.05x+0.25y=33 0.315x+0.05y=36 Find x and y. This question confuses me, kind of because all the not completely whole numbers throw me off. Please help!

OpenStudy (anonymous):

A way to do that is multiply the second equation by 5. You will have: 0.05x + 0.25y = 33 (i) 1,575x + 0,25y = 165 (ii) Now, do (ii) - (i) and you will have an equation only with the variable x. Do you get it?

OpenStudy (zzelinski):

Yes, thanks! I didn't think of multiplying the second equation by 5, it makes sense though!